OpenStudy (s):

If we approximate the atmosphere to be 84% Nitrogen and 16% Oxygen estimate the density of air (kg/m^3) at STP conditions (0c, 1atm).

OpenStudy (jamesj):

By the perfect gas law, \[ \frac{PV}{T} = nR \] where here \( n \) is the weighted sum of the nitrogen and oxygen. Let \( V = 1 \ m^3 \) and now you can figure out how many moles of O of N there must be in that space; hence the weight; hence the density.

OpenStudy (s):

Thank you but that's still not ringing a bell. This is for my environmental engineering class and I took chem and physics like 3 years ago so I can't remember much =\ Could you please tell me how to go on from here?

OpenStudy (s):

So iv V = 1, V of nitrogen is 0.84, and V of Oxygen is 0.16...

OpenStudy (jamesj):

Using T = 0 C = 273 K, V = 1 m^3 P = 1 atm R = 8.314 J/(K.mole) you can solve for n, the number of moles. Then in that space there are 0.84n moles of N 0.16n moles of O From that you can calculate the mass of the N and O, write them as \( m_O \) and \( m_N \). Thus the total mass is \( m = m_O + m_N \) and hence the density \[ \rho = \frac{m}{V} = \frac{(m_O + m_N) \ kg }{1 \ m^3} = (m_O + m_N) \ kg/m^3 \]

OpenStudy (jamesj):

(Notice in this last equation the m^3 is a measure of volume, not the mass cubed.)

OpenStudy (s):

Okay this makes perfect sense! Great explanation! Now i get it! Thank you very very much!